# E-SQL: Direct Schema Linking via Question Enrichment in Text-to-SQL

### v1

@@ 2.2.3. Data Augmentation
Data augmentation is a critical sub-task in Text-to-SQL, particularly in fine-tuning open-source LLMs using synthesized data. Li et al.  (Li et al., 2024b) underscore the importance of synthetic SQL-question generation and introduce a bi-directional augmentation technique to enhance model performance. Another essential aspect of data augmentation involves providing relevant content to the LLM through prompts. Augmenting prompts with items pertinent to the query is crucial for improving Text-to-SQL translation performance. Depending on the sub-task, prompts are typically enriched with explanations of database items, database values selected based on similarity, sub-task specific examples, and database schema—represented either as code or natural language—that are filtered or unfiltered, along with decomposed queries and candidate SQL queries (Gao et al., 2024; Pourreza and Rafiei, 2023; Talaei et al., 2024; Qu et al., 2024; Wang et al., 2024; Pourreza and Rafiei, 2024; Maamari et al., 2024). In our work, we enhance LLM performance by providing database item explanations, database values similar to the query, candidate conditions with exact database values, and possible execution errors for candidate queries to effectively bridge the gap between the natural language query and the database.
2.2.4. SQL Query Generation and Correction
@@
The final step in Text-to-SQL tasks is generating an SQL query that accurately answers the user’s natural language question. While some approaches, like C3 (Dong et al., 2023), employ zero-shot generation, few-shot prompting techniques are more commonly used to enhance performance. Methods such as self-consistency (Wang et al., 2023) and multi-choice selection are employed by models like C3, DAIL-SQL, CHESS, and MSC-SQL (Dong et al., 2023; Gao et al., 2024; Talaei et al., 2024; Lee et al., 2024; Maamari et al., 2024). To address missing or redundant keywords in generated SQLs, Pourreza and Rafiei (Pourreza and Rafiei, 2023) introduced a self-correction module with two versions: generic and gentle. The generic self-correction assumes that the initial SQL query contains errors and seeks to identify and correct them, while the gentle version detects potential issues and provides hints for improvement. With MAC-SQL, Wang et al. (Wang et al., 2024) proposed a refiner agent that evaluates the syntactic correctness and executability of the generated SQL, ensuring non-empty result generation; SQL correction is triggered if any issues are detected. In our work, we propose a SQL refinement module similar to MAC-SQL (Wang et al., 2024), where we refine the SQL after enriching the question and augmenting the prompt with candidate predicates.
@@ 3. Methodology
Our work is inspired by Self-Polish (Xi et al., 2023) and DBTagger (Usta et al., 2021), focusing on bridging the gap between the database schema and users’ natural language queries by enriching the questions with database items through keyword mappings and expanding them with SQL generation steps and reasoning. Our approach can be characterized as direct schema linking via query enrichment with relevant database elements. To further minimize errors such as incorrect table, column, or missing value usage in predicates, we augment the prompt with all possible predicates after extracting them.
Our proposed method consists of four main modules, as illustrated in Figure 2: Candidate SQL Generation (CSG), Candidate Predicate Generation (CPG), Question Enrichment (QE), and SQL Refinement (SR). Additionally, recognizing the positive impact of database schema filtering demonstrated in prior work, we include a Schema Filtering and Filtered Schema Correction module (SF). However, our experiments reveal that schema filtering can lead to performance degradation when using Text-to-SQL pipelines with advanced large language models (LLMs), corroborating the findings of Maamari et al. (Maamari et al., 2024). Detailed explanations of each module are provided in the following subsections, while the execution flow of our pipeline for a specific example is presented in Appendix C.
@@ 3.1. Database Description and Value Selection
Li et al. (Li et al., 2023b), Scholak et al. (Scholak et al., 2021), and Li et al. (Li et al., 2023c) utilize the Longest Common Substring (LCS) algorithm to identify database values related to the query. However, the time complexity of the LCS algorithm can significantly slow down the retrieval process. To address this issue, Li et al. (Li et al., 2024b) propose a ”coarse-to-fine” matching framework, which initially extracts hundreds of potentially relevant values from the entire database using BM25, followed by the application of the LCS algorithm to retrieve the most query-relevant values. Talaei et al. (Talaei et al., 2024) employ a different approach by first extracting keywords from the natural language query with the assistance of an LLM, and then retrieving similar values from the database using a Locality Sensitive Hashing (LSH) based hierarchical retrieval strategy. DIN-SQL (Pourreza and Rafiei, 2023) takes a more focused approach by extracting possible cell values directly from the query with the help of an LLM, similar to the keyword extraction phase of CHESS (Talaei et al., 2024). In our work, we employ the BM25 ranking algorithm not only to retrieve the most relevant database values but also to identify the most pertinent database descriptions. We augment prompts with the 10 most relevant database values for each column and the 20 most relevant database description sentences, as determined by the BM25 ranking algorithm. Additionally, when providing database values for each column, we ensure that any columns containing "NULL" values also include "NULL" as part of the selected values. This guarantees that the LLM is aware of potential null entries, allowing it to incorporate conditions such as "IS NOT NULL" when needed.
@@ 3.2. Candidate SQL Generation Module (CSG)
Our candidate SQL generation module incorporates three randomly selected samples, each from a different database than the one associated with the current query, across various difficulty levels from the few-shot data. These samples are ordered to present simpler question-SQL pairs first, followed by more challenging pairs. After providing database schema code representation as suggested in DAIL-SQL (Gao et al., 2024), the prompt is further augmented with selected database descriptions and relevant database values for each column as described in the previous subsection 3.1. To enhance the LLM’s reasoning capabilities, we use the phrase ”Let’s think step by step” and instruct the LLM to generate reasoning steps, as proposed by Kojima et al. (Kojima et al., 2022) and Wei et al. (Wei et al., 2024). The full prompt template for candidate SQL generation is provided in Appendix A.1.
@@ 3.3. Candidate Predicate Generation (CPG)
Here, <VALUE> represents the tokens extracted from the candidate SQL query. This process results in a list of possible predicates, formatted as ”<table>.<column> <operation> <value>,” which is used in downstream tasks.
@@
Figure 3. Example for the generation of incomplete value in the predicate explained in the Section 3.3, case (2).
@@
Figure 4. Example for the generation of incomplete value in the predicate explained in the Section 3.3, case (3).
@@
Figure 5. Example for the generation of incomplete value in the predicate explained in the Section 3.3, case (4).
@@ 3.4. Schema Filtering and Filtered Schema Correction Module (SF)
In prior works, schema filtering—eliminating database tables and columns irrelevant to the query—has been shown to improve model performance by reducing the likelihood of generating irrelevant schema items in SQL queries. One approach involves instructing the LLM to first select the relevant database tables and then choose the most relevant columns from those tables  (Dong et al., 2023; Talaei et al., 2024; Lee et al., 2024). Another approach filters the database schema by considering both tables and columns simultaneously in a single step  (Pourreza and Rafiei, 2023). Additionally, some methods leverage dummy SQL queries to extract relevant database entities for schema filtering  (Qu et al., 2024). Li et al. propose CODES  (Li et al., 2024b) and train a schema classifier to predict relevance scores following RESDSQL  (Li et al., 2023c). In our work, we adopt a single-step schema filtering approach, extracting only the relevant tables and their associated columns.
However, we observed that the filtered schema may not always be compatible with the original schema, where a selected column might belong to a different table according to the original database schema. When such issues arise, they can negatively impact the SQL generation process and lead to a decline in performance. To address this, we propose a filtered schema correction strategy, which checks for mismatches between the filtered schema and the original schema and corrects them accordingly. While previous research has demonstrated the benefits of schema filtering for schema linking and overall Text-to-SQL translation performance, our experiments show that schema filtering can decrease performance when used with the latest LLMs, a finding that aligns with the work of Maamari et al.  (Maamari et al., 2024). Detailed experimental results are provided in Section 4.3. Consequently, we have chosen not to include a schema filtering module in our E-SQL pipeline. The full prompt template used for Schema Filtering is provided in Appendix A.4
@@ 3.5. Question Enrichment Module (QE)
In the Question Enrichment module (QE), we instruct the LLM to refine the original question by incorporating relevant database items (tables, columns, and values) and conditions, making the question more understandable, clear, and free of irrelevant information. A few-shot strategy is applied to generate refined questions, using randomly sampled question-SQL pairs from the development set that have been manually annotated and made publicly available. The rationale for sampling from the development set is to ensure that we annotate 12 questions for each difficulty level: simple, moderate, and challenging. Since the manually annotated questions are taken from the development set, we ensure that the randomly selected 3 examples from each difficulty level provided in the question enrichment prompt are related to a database different from the one associated with the current query. This approach prevents enriched question examples that are directly related to the database of the considered query, ensuring that the query remains distinct in the enrichment process. In the few-shot examples, we include not only the enriched question but also the enrichment reasoning, manually prepared using the chain-of-thought technique, to fully leverage reasoning capabilities. The question enrichment prompt, whose template is provided in Appendix A.2, also includes the database schema, relevant database descriptions and values, and candidate predicates generated by the Candidate Predicate Generation (CPG) module. Examples of manually enriched questions, along with the rationale for their enrichment, are detailed in Appendix B.
@@
We observed that after applying the Question Enrichment module (QE), enriched questions may not always include the logical steps necessary to generate a correct SQL query, while the generated reasoning often contains these steps. To create fully enriched questions, we combine the original question, the enriched question generated by the LLM, and the enrichment reasoning generated by the LLM, as illustrated in 6(b). Unlike Self-Polish, we do not perform iterative question refinement, leaving iterative enrichment as a potential direction for future work.
@@ 3.6. Predicate and Error-Aware SQL Refinement Module (SR)
Since the generated SQL queries may contain minor mistakes, a common approach to address these is to use a correction module, as shown in Figure 1. Pourreza et al.  (Pourreza and Rafiei, 2023) proposed a novel self-correction module that relies heavily on LLMs to identify and correct potential errors without executing the SQL queries. Another approach employed by C3 (Dong et al., 2023), DAIL-SQL (Gao et al., 2024), CHESS (Talaei et al., 2024), and MCS-SQL (Lee et al., 2024) is to generate multiple SQL queries for a natural language question and select the most consistent one as the final predicted SQL, rather than relying solely on the greedily generated SQL. The refiner agent proposed by Wang et al.  (Wang et al., 2024) verifies the syntactic correctness and executability of the generated SQL, triggering a correction operation if the SQL fails these checks. The refiner agent performs these actions iteratively until the result is correct or the maximum number of iterations is reached. Although this design enhances performance, it also increases both the cost and time required to produce the final SQL query. In our work, as illustrated in Figure 2, we execute the candidate SQL query and detect any execution errors. Using this error information, along with candidate predicates and the enriched question, we instruct the LLM to either generate a new SQL query or refine the existing candidate SQL query by incorporating the question context, database schema, candidate predicates, and execution error details. The full prompt template used for SQL Refinement is provided in Appendix A.3.
4. Experiments
@@ 4.1.1. Dataset
We conduct our experiments on the most challenging cross-domain BIRD dataset  (Li et al., 2024a). BIRD focuses on the real-word complex and large databases, and it provides external knowledge to enhance the capability of large language models to understand both the question and the database structure and values better. BIRD dataset spans 37 professional domains such as football, formula 1, blockchain, healthcare, and education, etc., and it contains 12,751 text-to-SQL pairs from 95 databases with a size of 33.4 GB. The training set consists of 9,428 text-to-SQL pairs, while the development and test sets consist of 1,534 and 1,789 instances, respectively. The test set of the BIRD dataset is concealed.
@@ 4.1.3. Models
In this work, we employed the latest proprietary models, GPT-4o-mini and GPT-4o, as the backbone for our experiments. Exploring other proprietary models or fine-tuning open-source LLMs are left as future work. The majority of the experiments were conducted using GPT-4o-mini, which is approximately 30% more cost-effective than GPT-4o. Despite its lower cost, GPT-4o-mini demonstrated robust capabilities, making it an excellent choice for balancing performance and budget constraints.
@@ 4.2. Results
We evaluate the proposed E-SQL pipeline on the BIRD development set since the test set is not publicly available. Since the GPT-4o-mini is more cost-effective, we conduct most of our experiments using GPT-4o-mini. Table 1 compares the performance of E-SQL and several baseline models, demonstrating competitive results. Additionally, we evaluate the performance of our method across various difficulty levels.
E-SQL shows a notable improvement, particularly for challenging questions, largely due to the incorporation of the Question Enrichment (QE) and SQL Refinement (SR) modules. With GPT-4o, the pipeline achieved an Execution Accuracy (EX) of 66.29 and a Soft F1-Score of 67.93 on the test set. When using the more cost-effective model, GPT-4o-mini, E-SQL achieved 59.81 EX and 61.59 Soft F1-Score on the test set. The detailed breakdown of performance across different query complexity levels on the BIRD test set can be found in Table 2. These results highlight the efficacy of question enrichment, candidate predicate augmentation, and schema refinement techniques, especially in handling complex queries.
@@
Method dev EX test EX R-VES
AskData 72.03 72.39 66.92
@@
Table 2. Detailed Performance of E-SQL on BIRD Test Set Using GPT-4o and GPT-4o-mini Across Query Complexity Levels
@@
Table 4. Ablation study using GPT-4o-mini with EX and Soft F1 metrics on the development set. The arrows indicate performance improvement ($\uparrow$) or decline ($\downarrow$) compared to the base E-SQL.

### v2

@@ 2.3. Data Augmentation
One essential aspect of data augmentation involves providing relevant content to the LLM through prompts. Augmenting prompts with items pertinent to the query is crucial for improving Text-to-SQL translation performance. Depending on the sub-task, prompts are typically enriched with explanations of database items, database values selected based on similarity, sub-task specific examples, and database schema—represented either as code or natural language—that are filtered or unfiltered, along with decomposed queries and candidate SQL queries (Gao et al., 2024; Pourreza and Rafiei, 2023; Talaei et al., 2024; Qu et al., 2024; Wang et al., 2024; Pourreza and Rafiei, 2024; Maamari et al., 2024). In our work, we incorporate database item explanations and database values similar to the query in prompt. Additionally, we include execution errors for candidate queries to provide valuable feedback during query refinement. Unlike prior work, we propose a novel approach to further reduce predicate errors and enhance LLM performance by introducing candidate conditions with exact database values, effectively bridging the gap between the natural language query and the database.
2.4. SQL Query Generation and Correction
@@
The final step in Text-to-SQL tasks is generating an SQL query that accurately answers the user’s natural language question. While some approaches, like C3 (Dong et al., 2023), employ zero-shot generation, few-shot prompting techniques are more commonly used to enhance performance. Methods such as self-consistency (Wang et al., 2023) and multi-choice selection are employed by models like C3, DAIL-SQL, CHESS, and MSC-SQL (Dong et al., 2023; Gao et al., 2024; Talaei et al., 2024; Lee et al., 2024; Maamari et al., 2024). However, these techniques lead to high computational costs due to large number of generation steps for a single query. To address missing or redundant keywords in generated SQLs, Pourreza and Rafiei (Pourreza and Rafiei, 2023) introduced a self-correction module. With MAC-SQL, Wang et al. (Wang et al., 2024) proposed a refiner agent that evaluates the syntactic correctness and executability of the generated SQL, ensuring non-empty result generation. In our work, we propose a SQL refinement module similar to MAC-SQL (Wang et al., 2024), where we refine the SQL after enriching the question and augmenting the prompt with candidate predicates.
@@ 3. Methodology
Our work focuses on bridging the gap between the database schema and users’ natural language queries by enriching the questions with database items through keyword mappings and expanding them with SQL generation steps and reasoning. Our approach can be characterized as direct schema linking via query enrichment with relevant database elements. To further minimize errors such as incorrect table, column, or missing value usage in predicates, we augment the prompt with all possible predicates after extracting them.
Our proposed method consists of four main modules, as illustrated in Figure 2: Candidate SQL Generation (CSG), Candidate Predicate Generation (CPG), Question Enrichment (QE), and SQL Refinement (SR). Additionally, recognizing the mixed outcomes of database schema filtering observed in prior work, we include a Schema Filtering and Filtered Schema Correction module (SF). However, our experiments on our pipeline reveal that schema filtering can lead to performance degradation when integrated into Text-to-SQL pipelines with the most advanced large language models (LLMs), corroborating the findings of Maamari et al. (Maamari et al., 2024). Detailed explanations of each module are provided in the following subsections.
@@ 3.1. Database Description and Value Selection
Li et al. (Li et al., 2023b), Scholak et al. (Scholak et al., 2021), and Li et al. (Li et al., 2023c) utilize the Longest Common Substring (LCS) algorithm to identify database values related to the query. However, the time complexity of the LCS algorithm can significantly slow down the retrieval process. To address this issue, Li et al. (Li et al., 2024c) propose a ”coarse-to-fine” matching framework, which initially extracts hundreds of potentially relevant values from the entire database using BM25, followed by the application of the LCS algorithm to retrieve the most query-relevant values. Talaei et al. (Talaei et al., 2024) employ a different approach by first extracting keywords from the natural language query with the assistance of an LLM, and then retrieving similar values from the database using a Locality Sensitive Hashing (LSH) based hierarchical retrieval strategy. DIN-SQL (Pourreza and Rafiei, 2023) takes a more focused approach by extracting possible cell values directly from the query with the help of an LLM, similar to the keyword extraction phase of CHESS (Talaei et al., 2024). In our work, we employ the BM25 ranking algorithm not only to retrieve the most relevant database values but also to identify the most pertinent database descriptions. We augment prompts with the 10 most relevant database values for each column and the 20 most relevant database description sentences, as determined by the BM25 ranking algorithm. Additionally, when providing database values for each column, we ensure that any columns containing "NULL" values also include "NULL" as part of the selected values. This guarantees that the LLM is aware of potential null entries, allowing it to incorporate conditions such as "IS NOT NULL" when needed.
@@ 3.2. Candidate SQL Generation Module (CSG)
Our candidate SQL generation module incorporates three randomly selected samples, each from a different database than the one associated with the current query, across various difficulty levels from the few-shot data. These samples are ordered to present simpler question-SQL pairs first, followed by more challenging pairs. After providing database schema code representation as suggested in DAIL-SQL (Gao et al., 2024), the prompt is further augmented with selected database descriptions and relevant database values for each column as described in the previous subsection 3.1. To enhance the LLM’s reasoning capabilities, we use the phrase ”Let’s think step by step” and instruct the LLM to generate reasoning steps, as proposed by Kojima et al. (Kojima et al., 2022) and Wei et al. (Wei et al., 2024).
@@ 3.3. Candidate Predicate Generation (CPG)
Here, <VALUE> represents a token extracted from the values in candidate SQL query. This process results in a list of possible predicates, formatted as ”<table>.<column> <operation> <value>,” which is used in downstream tasks.
@@
Figure 3. Example for the generation of incomplete value in the predicate explained in Section 3.3, case (2).
@@
Figure 4. Example for correct table and value but wrong column in the predicate explained the Section 3.3, case (3).
@@
Figure 5. Example for correct table but wrong column and value selection in the predicate explained in Section 3.3, case (4).
@@ 3.4. Schema Filtering and Filtered Schema Correction Module (SF)
In prior works, schema filtering—eliminating database tables and columns irrelevant to the query—has been shown to improve model performance by reducing the likelihood of generating irrelevant schema items in SQL queries. One approach involves instructing the LLM to first select the relevant database tables and then choose the most relevant columns from those tables  (Dong et al., 2023; Talaei et al., 2024; Lee et al., 2024). Another approach filters the database schema by considering both tables and columns simultaneously in a single step  (Pourreza and Rafiei, 2023). Additionally, some methods leverage dummy SQL queries to extract relevant database entities for schema filtering  (Qu et al., 2024). Li et al. propose CODES  (Li et al., 2024c) and train a schema classifier to predict relevance scores following RESDSQL  (Li et al., 2023c). In our work, we adopt a single-step schema filtering approach, extracting only the relevant tables and their associated columns.
However, we observed that the filtered schema may not always be compatible with the original schema, where a selected column might belong to a different table according to the original database schema. When such issues arise, they can negatively impact the SQL generation process and lead to a decline in performance. To address this, we propose a filtered schema correction strategy, which checks for mismatches between the filtered schema and the original schema, subsequently correcting them accordingly. While previous research has demonstrated the benefits of schema filtering for schema linking and overall Text-to-SQL translation performance, our experiments show that incorporation of schema filtering into our pipeline result in performance decrease when used with the latest proprietary LLMs, a finding that aligns with the work of Maamari et al.  (Maamari et al., 2024). Detailed experimental results are provided in Section 4.3. Consequently, we have chosen not to include a schema filtering module in our E-SQL pipeline.
@@ 3.5. Question Enrichment Module (QE)
In the Question Enrichment module (QE), we instruct the LLM to refine the original question by incorporating relevant database items (tables, columns, and values) and conditions. This process makes the question more understandable, coherent, well-aligned with the database schema. Additionally, an SQL construction plan, generated during the question-database integration as part of the reasoning process, is appended to the enriched question. Together, this plan and the enriched question form the fully enriched question which explicitly incorporates the necessary SQL components and logical steps, guiding the LLM to generate accurate SQL queries. The creation process of a fully enriched question, which combines the original question, the enriched question, and the enrichment reasoning, is illustrated in  6(b). A few-shot strategy is applied to generate refined questions, using randomly sampled question-SQL pairs from the development set that have been manually annotated and made publicly available. We annotate 12 questions for each difficulty level: simple, moderate, and challenging. Since the manually annotated questions are taken from the development set, we ensure that the randomly selected 3 examples from each difficulty level provided in the question enrichment prompt are related to a database different from the one associated with the current query. This approach prevents enriched question examples directly related to the database of the considered query from being included in the prompt. In the few-shot examples, we include both the enriched question and the enrichment reasoning, manually prepared considering the chain-of-thought technique, to fully leverage reasoning capabilities of the model and explicitly outline the SQL logical steps. The question enrichment prompt also includes the database schema, relevant database descriptions and values, and candidate predicates generated by the Candidate Predicate Generation (CPG) module. The fully enriched question is generated in a single prompt, without iterative refinement. Iterative enrichment is left as a potential direction for future work.
@@ 3.6. Predicate and Error-Aware SQL Refinement Module (SR)
Since the generated SQL queries may contain minor mistakes, a common approach to address these is to use a correction module, as shown in Figure 1. Pourreza et al.  (Pourreza and Rafiei, 2023) proposed a novel self-correction module that relies heavily on LLMs to identify and correct potential errors without executing the SQL queries. Another approach employed by C3 (Dong et al., 2023), DAIL-SQL (Gao et al., 2024), CHESS (Talaei et al., 2024), and MCS-SQL (Lee et al., 2024) is to generate multiple SQL queries for a natural language question and select the most consistent one as the final predicted SQL, rather than relying solely on the greedily generated SQL. The refiner agent proposed by Wang et al.  (Wang et al., 2024) verifies the syntactic correctness and executability of the generated SQL, triggering a correction operation if the SQL fails these checks. The refiner agent performs these actions iteratively until the result is correct or the maximum number of iterations is reached. Although this design enhances performance, it also increases both the cost and time required to produce the final SQL query. In our work, as illustrated in Figure 2, we execute the candidate SQL query and detect any execution errors. Using this error information, along with candidate predicates, the enriched question, and database schema, we instruct the LLM to either generate a new SQL query or refine the existing candidate SQL query.
The prompt skeletons used for each module in our pipeline are provided in our Github repository. Furthermore, a detailed example illustrating the complete workflow of the pipeline for a sample user query is also provided for reference.
4. Experiments and Results
@@ 4.1.1. Dataset
We conduct our experiments on the Spider (Yu et al., 2018) dataset and the most challenging cross-domain BIRD dataset  (Li et al., 2024a). Spider benchmark includes 10,181 Text-to-SQL pairs from 200 databases spanning 138 domains, each containing multiple tables. For our experiments, we used the test split of the Spider dataset, which consists of 2,147 Text-to-SQL pairs. BIRD focuses on the real-word complex and large databases, and it provides external knowledge to enhance the capability of large language models to understand both the question and the database structure and values better. BIRD dataset spans 37 professional domains such as football, formula 1, blockchain, healthcare, and education, etc., and it contains 12,751 text-to-SQL pairs from 95 databases with a size of 33.4 GB. The training set consists of 9,428 Text-to-SQL pairs, while the development and test sets consist of 1,534 and 1,789 instances, respectively. The test set of the BIRD dataset is concealed.
@@ 4.1.3. Models
In this work, we employed the latest proprietary models, (GPT-4o-mini and GPT-4o) and small open-source LLMs (Qwen2.5 Coder (Hui et al., 2024), DeepSeek Coder (Guo et al., 2024)) without fine-tuning as the backbone for our experiments. Exploring other models or fine-tuning open-source LLMs are left as future work. For proprietary models, the majority of the experiments were conducted using GPT-4o-mini, which is approximately 30% more cost-effective than GPT-4o. Despite its lower cost, GPT-4o-mini demonstrated robust capabilities, making it an excellent choice for balancing performance and budget constraints.
@@ 4.2. Results
We evaluate the proposed E-SQL pipeline primarily on the Spider test set and the BIRD development set, as the BIRD test set is not publicly available. We conduct most of our experiments using GPT-4o-mini due to its cost-effectiveness. We also experiment with small open-source models without fine-tuning to evaluate the impact of the pipeline and highlight the importance of database-aligned questions. Table 1 compares the performance of E-SQL and several baseline models on the BIRD dataset, demonstrating competitive results. Additionally, we evaluate the performance of our method across various difficulty levels on BIRD dataset. Table 2 provides insights into the performance of E-SQL utilizing cost-effective proprietary LLMs and small open-source LLMs on the Spider test dataset.
E-SQL shows a notable improvement, particularly for challenging questions in BIRD dataset, largely due to the incorporation of the Question Enrichment (QE) and SQL Refinement (SR) modules. With GPT-4o, the pipeline achieved an Execution Accuracy (EX) of 66.29 and a Soft F1-Score of 67.93 on the BIRD test set. When using the more cost-effective model, GPT-4o-mini, E-SQL achieved 59.81 EX and 61.59 Soft F1-Score on the BIRD test set. Additionally, using the small open-source LLM Qwen2.5 Coder 7B, E-SQL achieved 53.59 EX on the BIRD development set. On the Spider test set, E-SQL achieved 74.75 EX score using GPT-4o-mini and 58.64 EX score using Qwen2.5 Coder 7B Instruct. The detailed breakdown of performance across different datasets and query complexity levels can be found in Table 3 and Table 6. Further insights into the impact of question enrichment and candidate predicate augmentation are provided in Sections 4.4 and 4.5. These results highlight the efficacy of question enrichment, candidate predicate augmentation, and schema refinement techniques, especially in handling complex queries.
@@
Method Dev EX Test EX Test R-VES
Undisclosed
@@
Published
CHESS 65.00 66.69 62.77
MCS-SQL + GPT-4  (Lee et al., 2024) 63.36 65.45 61.23
SuperSQL  (Li et al., 2024b) 58.50 62.66 -
SFT CodeS-15B  (Li et al., 2024c) 58.47 60.37 61.37
DTS-SQL + DeepSeek 7B  (Pourreza and Rafiei, 2024) 55.80 60.31 -
MAC-SQL + GPT-4  (Wang et al., 2024) 57.56 59.59 57.60
SFT CodeS-7B  (Li et al., 2024c) 57.17 59.25 55.69
TA-SQL + GPT-4  (Qu et al., 2024) 56.19 59.14 -
DAIL-SQL + GPT-4  (Gao et al., 2024) 54.76 57.41 54.02
@@
E-SQL + Qwen2.5 Coder 7B Instruct (Ours) 53.59 - -
Table 2. Performance of E-SQL on Spider test set. E-SQL consists of CSG, CPG, QE, and SR modules as illustrated in Figure 2. symbol denotes methods utilizing fine-tuned LLMs.
Method Model Test EX
Proprietary Models
DAIL-SQL (Gao et al., 2024) GPT-4 86.6
DIN-SQL (Pourreza and Rafiei, 2023) GPT-4 85.3
TA-SQL (Qu et al., 2024) GPT-4 85.0
MAC-SQL (Wang et al., 2024) GPT-3.5-Turbo 75.5
- GPT4 74.0
E-SQL (ours) GPT-4o-mini 74.75
Small Open Source Models
DTS-SQL (Pourreza and Rafiei, 2024) Mistral-7B 77.1
MSc-SQL (Gorti et al., 2024) Gemma-2-9B 69.30
E-SQL (ours) Qwen2.5 Coder 7B Instruct 58.64
Table 3. Detailed Performance of E-SQL on BIRD Test Set Across Query Complexity Levels
@@
Table 5. Ablation study using GPT-4o-mini with EX and Soft F1 metrics on the BIRD development set. The arrows indicate performance improvement ($\uparrow$) or decline ($\downarrow$) compared to the base E-SQL.